# Capturing magic angles in twisted bilayer graphene from information theory markers

### v1

@@ Abstract
Zero energy eigenstates $\psi_{0}(\theta)$ of the twisted bilayer graphene Hamiltonian at the Dirac point show a high sensitivity to the twist angle $\theta$ near the magic angles where the effective Fermi velocity vanishes. We use information theory markers, like fidelity-susceptibility and entanglement entropy of the reduced density matrix to the layer sector, to capture this quantum criticality of zero modes at magic twist angles.
@@ I Introduction
After its rediscovery, isolation and investigation by Novoselov and Geim [1, 2], graphene and other 2D (thin layer) atomic crystals have become valuable and useful nanomaterials (see e.g. [3] for a review on the electronic properties of graphene). The theoretical identification of graphene as a 2D topological (quantum spin Hall) insulator by [4] also opened a field of interesting technological applications like spintronic, magnetoelectronic and optoelectronic devices and dissipationless transistors (see [5, 6] for traditional reviews).
@@
Quantum information science viewpoints are increasingly being introduced into condensed matter physics. Actually, the study of topological order and the related new quantum phases is actually a study of patterns of entanglement [26, 27]. Rényi-Wherl entropies and inverse participation ratio localization measures have proven to identify topological phase transitions in silicene [28, 29], phosphorene [30, 31] and HgTe/CdTe quantum wells [32, 33]. The concept of fidelity-susceptibility has been quite useful in describing traditional quantum phase transitions [34, 35, 36, 37, 38, 39, 40]. Here we want to explore this information tool in capturing some criticality at magic angles. Entanglement spectrum and lattice-layer entanglement has also been studied in [41] and [42] for the particular case of Bernal-stacked bilayer graphene. Here we extend these studies to the more general TBG case, recovering the previous as a particular case, and demonstrating how magic angle signatures are printed in these quantities.
@@ II.1 Momentum space, interlayer hopping and Hilbert space basis
where $s=A,B$ makes reference to the two triangular graphene sublattices and we just consider one single value of the spin third component (either up or down); we are measuring momentum $\bm{q}$ from the Dirac point $\bm{K}$ (similar results are obtained at $\bm{K}^{\prime}=-\bm{K}$). These states form an orthonormal basis $\langle\ell,\bm{m},s|\ell^{\prime},\bm{m}^{\prime},s^{\prime}\rangle=\delta_{\bm{m}\bm{m}^{\prime}}\delta_{\ell\ell^{\prime}}\delta_{ss^{\prime}}$, where the Kronecker $\delta_{\bm{m}\bm{m}^{\prime}}$ understood on a component-by-component basis.
@@
The quantity $({|\bm{q}_{\bm{m}}^{1}|}/{K_{\theta}})^{2}/3=m_{1}^{2}-m_{1}m_{2}+m_{2}^{2}$ appears in the literature as Löschian numbers (the norm of the Eisenstein integers), with interesting properties in number theory. Note that $m_{1}^{2}-m_{1}m_{2}+m_{2}^{2}=n$ defines ellipses in the $(m_{1},m_{2})$ plane (see Fig. 2, right panel). In Figure 2 (left panel) we draw wave vectors $\bm{q}_{\bm{m}}^{\ell}$ in $K_{\theta}$ units as blue dots for layer $\ell=1$ and red dots for layer $\ell=2$. This allows us to arrange the spinor basis states $|\ell,\bm{m},s\rangle$ according to the distance $R_{\ell}={|\bm{q}_{\bm{m}}^{\ell}|}/{K_{\theta}}$ to the origin of the Moiré reciprocal lattice. The Hamiltonian matrix size $D$ coincides with twice (due to the sublattice $s$ degree of freedom) the number of dots enclosed inside the circle of radius $R_{\ell}$. The minimum Hamiltonian matrix size corresponding to the first three NN of layer $\ell=2$, $(m_{1},m_{2})=(0,0),(0,1)$ and $(-1,0)$ (represented by blue dots on the smallest blue circle of radius $R_{2}=1$ in Fig. 2) is $D=8$; an explicit expression of the matrix Hamiltonian for this case is given in [10]). For ten Moiré reciprocal lattice vectors (represented by 3 red and 7 blue dots inside the smallest blue circle of radius $R_{1}=\sqrt{3}$ in Fig. 2) one obtains a Hamiltonian matrix size $D=20$, and so on. In general, the Hilbert space dimension $D$ is roughly proportional to the area of the circle of radius $R$. A linear fit of the first 109 Moiré reciprocal lattice vectors inside a radius $R=\sqrt{43}$ gives
@@
Figure 2: Left panel: wave vectors $\bm{q}_{\bm{m}}^{\ell}$ of Eq. (2) (in $K_{\theta}$ units) are plotted as blue (layer $\ell=1$) and red (layer $\ell=2$) dots. We arrange them in circles of radius $R=|\bm{q}_{\bm{m}}^{\ell}|/K_{\theta}=1,\sqrt{3},2,\sqrt{7},3,\dots,$ which yields Hamiltonian matrix orders of: $D=8,20,26,38,50,\dots$. Right panel: ellipses ${|\bm{q}_{\bm{m}}^{\ell}|}/{K_{\theta}}=\sqrt{43}$ are drawn in the plane $(m_{1},m_{2})$; common points to both ellipses $\ell=1,2$ are plotted in black color.
@@ II.2 Hamiltonian matrix and electronic spectrum
$\mathcal{T}_{\bm{q}_{tl}}=w\begin{pmatrix}e^{-i\phi}&1\\ e^{-i\phi}&e^{-i\phi}\end{pmatrix},\;\mathcal{T}_{\bm{q}_{tr}}=w\begin{pmatrix}e^{i\phi}&1\\ e^{-i\phi}&e^{i\phi}\end{pmatrix},$
@@ III Effective Fermi velocity, flat bands, zero modes and magic angles
There is an “anomalous” pick of $v_{*}$ at $\theta\simeq 0.4^{\circ}$, as can be seen in Fig. 4, which also remains conspicuous for other quantum information markers (fidelity) explained in the next section. This is related to a degeneracy $E_{0}^{\pm}\simeq E_{1}^{\pm}$ between the lowest (quasi-zero) energy $\psi_{0}^{\pm}$ and the first excited $\psi_{1}^{\pm}$ Hamiltonian eigenstates at $\theta\simeq 0.4^{\circ}$, as can be seen in Fig. 5. We have also revealed the existence of “excited magic angles” like $\theta\simeq 0.52,0.66^{\circ}$ where the Fermi velocity expectation value $|\langle\psi_{1}^{\pm}|\bm{V}_{*}|\psi_{1}^{\pm}\rangle|$ in the first excited states $\psi_{1}^{\pm}$ vanishes. The possible relevance and physical significance of these excited magic angles is yet to be determined.
@@
Figure 5: Energies of quasi-zero $\psi_{0}^{\pm}$ (thickest curves) and first excited $\psi_{1}^{\pm}$ Hamiltonian eigenstates [conduction (+) in red and valence (-) in blue] as a function of the twist angle $\theta$ (in degrees). Energies $E_{0}^{\pm}$ and $E_{1}^{\pm}$ degenerate at the twist angle $\theta=0.4^{\circ}$ (marked with a grey vertical gridline). This degeneracy seems to be responsible for the sudden changes of the Fermi velocity and fidelity measures between the magic angles $\theta_{2}$ and $\theta_{3}$ seen in Fig. 4.
@@ IV.1 Fidelity-susceptibility
In Figure 4 (middle panel) we plot the fidelity $F_{\psi_{0}^{\pm}}$ of the lowest energy (quasi-zero) $E_{0}^{\pm}$ Hamiltonian eigenstates $\psi_{0}^{\pm}$ at the Dirac point $\bm{q}=\bm{0}$ as a function of $\theta\in[0.18,1.3]$ in degrees (we take a step $\delta\theta=0.0023^{\circ}$). We identify sudden drops of fidelity exactly at the magic angles $\theta_{1}=1.05^{\circ}$ (principal), $\theta_{2}=0.495^{\circ},\theta_{3}=0.35^{\circ},\theta_{4}=0.235^{\circ}$ and $\theta_{5}=0.193^{\circ}$ where the effective Fermi velocity $v_{*}$ drops to zero. Additionally, fidelity detects the twist angle $\theta\simeq 0.4^{\circ}$ where $\psi_{0}^{\pm}$ degenerates with $\psi_{1}^{\pm}$ (see Fig. 5). In Figure 4 (bottom panel) we see that magic angles are also detected by sudden rises of the overlap
@@ IV.2.2 RDM and entanglement at magic angles
Figure 8: Linear entropy and its derivative for the lowest-energy valence (similar results for conduction) eigenstate $\psi_{0}^{-}$ as a function of the twist angle.
@@ V Conclusions
We have contributed to unraveling the magic behind the electronic properties and exotic quantum states of TBG, which make it a promising platform for future device applications in electronics, optoelectronics, and quantum computing.
@@
We have reviewed the flattening of low-energy Moiré Bloch bands and the formation of sharp picks in the DOS at magic angles $\theta\simeq 1.05^{\circ},0.495^{\circ},0.35^{\circ},0.235^{\circ}$ and $0.193^{\circ}$. This flattening is accompanied by a vanishing of the effective Dirac-point Fermi velocity for zero energy Hamiltonian eigenstates $\psi_{0}(\theta)$ at those magic angles. Zero modes undergo critical structural changes at magic angles, and these abrupt changes are clearly detected by fidelity measurements. Fidelity also captures an anomalous behavior that the Fermi velocity already adverted at $\theta\simeq 0.4^{\circ}$, where the energies of the zero modes $\psi_{0}(\theta)$ and first excited $\psi_{1}(\theta)$ states degenerate. Sharp oscillations (like “resonances”) at magic angles are also noticeable in the entanglement entropy and the interlayer hopping probabilities (off-diagonal component of the RDM to the layer sector) as a function of the the twist angle. These resonances are better detected by the corresponding derivatives. Entanglement entropy is almost maximum in the small angle region where magic angles fall, and it decreases to zero for larger values of $\theta$. Low-energy electrons seem to prefer one layer over the other for large $\theta$, as seen in the layer probabilities $P_{\ell}$, which oscillates around $P_{\ell}=0.5$ in the small angle region where magic angles fall. The interlayer hopping probability $T$ shows high variability for small angles and goes to zero for larger values of $\theta$.
@@
Tunability with $\theta$ and manipulation with external electric and magnetic fields in general 2D materials allows for the exploration of novel quantum phenomena termed as “twistronic” [50, 8]. We believe that quantum information viewpoints can contribute to a better understanding for the physics behind and to the design of new experiments. Novel quantum phenomena and new technological possibilities are also expected when adding a stacking degree of freedom and going from bilayer to multilayer arrangements. For example, twisted trilayer graphene seems to provide a promising platform for experimental realization of correlated topological states [51, 52]. Multilayer arrangements also increase the number of internal quantum degrees of freedom to the electron necessary to encode logical qubits. Therefore, investigations in this direction are worth to explore.

### v2

@@ Abstract
Abstract. Zero energy eigenstates $\psi_{0}(\theta)$ of the twisted bilayer graphene Hamiltonian at the Dirac point show a high sensitivity to the twist angle $\theta$ near the magic angles where the effective Fermi velocity vanishes. We use information theory markers, like fidelity-susceptibility and entanglement entropy of the reduced density matrix to the layer sector, to capture this quantum criticality of zero modes at magic twist angles.
@@ I Introduction
After its rediscovery, isolation and investigation by Novoselov and Geim [1, 2], graphene and other 2D (thin layer) atomic crystals have become valuable and useful nanomaterials (see e.g. [3] for a review on the electronic properties of graphene). The theoretical identification of graphene as a 2D topological (quantum spin Hall) insulator by [4] also opened a field of interesting technological applications like spintronic, magnetoelectronic and optoelectronic devices and dissipationless transistors (see [5, 6] for traditional reviews), like FETs based on Bernal stacked bilayer graphene [7].
@@
Quantum information science viewpoints are increasingly being introduced into condensed matter physics. Actually, the study of topological order and the related new quantum phases is actually a study of patterns of entanglement [27, 28]. More and more general methods to measure entanglement entropy in a many-body systems are proposed in the literature (see e.g. [29] for one of them based on a quantum switch coupled to several copies of the original system). Rényi-Wherl entropies and inverse participation ratio localization measures have proven to identify topological phase transitions in silicene [30, 31], phosphorene [32, 33] and HgTe/CdTe quantum wells [34, 35]. The concept of fidelity-susceptibility has been quite useful in describing traditional quantum phase transitions [36, 37, 38, 39, 40, 41, 42]. Here we want to explore this information tool in capturing some criticality at magic angles. Entanglement spectrum and lattice-layer entanglement has also been studied in [43] and [44] for the particular case of Bernal-stacked bilayer graphene. Here we extend these studies to the more general TBG case, recovering the previous as a particular case, and demonstrating how magic angle signatures are printed in these quantities.
@@ II.1 Momentum space, interlayer hopping and Hilbert space basis
where $s=A,B$ makes reference to the two triangular graphene sublattices and we just consider one single value of the spin third component (either up or down); we are measuring momentum $\bm{q}$ from the Dirac point $\bm{K}$ (similar results are obtained at $\bm{K}^{\prime}=-\bm{K}$). These states form an orthonormal basis $\langle\ell,\bm{m},s|\ell^{\prime},\bm{m}^{\prime},s^{\prime}\rangle=\delta_{\bm{m}\bm{m}^{\prime}}\delta_{\ell\ell^{\prime}}\delta_{ss^{\prime}}$, where the Kronecker $\delta_{\bm{m}\bm{m}^{\prime}}$ is understood on a component-by-component basis.
@@
The quantity $({|\bm{q}_{\bm{m}}^{1}|}/{K_{\theta}})^{2}/3=m_{1}^{2}-m_{1}m_{2}+m_{2}^{2}$ appears in the literature as Löschian numbers (the norm of the Eisenstein integers), with interesting properties in number theory. Note that $m_{1}^{2}-m_{1}m_{2}+m_{2}^{2}=n$ defines ellipses in the $(m_{1},m_{2})$ plane (see Fig. 2, right panel). In Figure 2 (left panel) we draw wave vectors $\bm{q}_{\bm{m}}^{\ell}$ in $K_{\theta}$ units as blue dots for layer $\ell=1$ and red dots for layer $\ell=2$. This allows us to arrange the spinor basis states $|\ell,\bm{m},s\rangle$ according to the distance $R_{\ell}={|\bm{q}_{\bm{m}}^{\ell}|}/{K_{\theta}}$ to the origin of the moiré reciprocal lattice. The Hamiltonian matrix size $D$ coincides with twice (due to the sublattice $s$ degree of freedom) the number of dots enclosed inside the circle of radius $R_{\ell}$. The minimum Hamiltonian matrix size corresponding to the first three NN of layer $\ell=2$, $(m_{1},m_{2})=(0,0),(0,1)$ and $(-1,0)$ (represented by red dots on the smallest red circle of radius $R_{2}=1$ in Fig. 2) is $D=8$; an explicit expression of the matrix Hamiltonian for this case is given in [11]. For ten moiré reciprocal lattice vectors (represented by 3 red and 7 blue dots inside the smallest blue circle of radius $R_{1}=\sqrt{3}$ in Fig. 2) one obtains a Hamiltonian matrix size $D=20$, and so on. In general, the Hilbert space dimension $D$ is roughly proportional to the area of the circle of radius $R$. A linear fit of the first 109 moiré reciprocal lattice vectors inside a radius $R=\sqrt{43}$ gives
@@
Figure 2: Left panel: wave vectors $\bm{q}_{\bm{m}}^{\ell}$ of Eq. (2) (in $K_{\theta}$ units) are plotted as blue (layer $\ell=1$) and red (layer $\ell=2$) dots. We arrange them in circles of radius $R=|\bm{q}_{\bm{m}}^{\ell}|/K_{\theta}=1,\sqrt{3},2,\sqrt{7},3,\dots,$ which yields Hamiltonian matrix orders of: $D=8,20,26,38,50,\dots$ Right panel: ellipses ${|\bm{q}_{\bm{m}}^{\ell}|}/{K_{\theta}}=\sqrt{43}$ are drawn in the plane $(m_{1},m_{2})$; common points to both ellipses $\ell=1,2$ are plotted in black color.
@@ II.2 Hamiltonian matrix and electronic spectrum
$\mathcal{T}_{\bm{q}_{tl}}=w\begin{pmatrix}e^{-\mathrm{i}\phi}&1\\ e^{\mathrm{i}\phi}&e^{-\mathrm{i}\phi}\end{pmatrix},\;\mathcal{T}_{\bm{q}_{tr}}=w\begin{pmatrix}e^{\mathrm{i}\phi}&1\\ e^{-\mathrm{i}\phi}&e^{\mathrm{i}\phi}\end{pmatrix},$
@@ III Effective Fermi velocity, flat bands, zero modes and magic angles
There is an outlier value of the effective Fermi velocity at the twist angle $\theta\simeq 0.4^{\circ}$, between the magic angles $\theta_{2}$ and $\theta_{3}$ in the top panel of Fig. 4, perceived as an anomalous pick when joining the points of the list plot. This outlier value persists for other Hamiltonian model parameters, like $v$ or $w$, and matrix sizes $D$, although its location varies with $v$ and $w$, so do the magic angles. Therefore, an explanation is necessary for this outlier. It turns out to be related to a degeneracy $E_{0}^{\pm}\simeq E_{1}^{\pm}$ between the lowest (quasi-zero) energy $\psi_{0}^{\pm}$ and the first excited $\psi_{1}^{\pm}$ Hamiltonian eigenstates at $\theta\simeq 0.4^{\circ}$, as can be seen in Fig. 5. Moreover, $\theta\simeq 0.4^{\circ}$ turns out to be an “excited magic angle" in the sense that the effective Fermi velocity expectation value
$v_{*}^{1}/v=|\langle\psi_{1}^{\pm}|\bm{V}_{*}|\psi_{1}^{\pm}\rangle|/v$ (13)
in the first excited states $\psi_{1}^{\pm}$ vanishes (see the inset in Fig 5). This degeneracy between $\psi_{0}^{\pm}$ and $\psi_{1}^{\pm}$ at $\theta\simeq 0.4^{\circ}$ also manifest itself in a sudden fidelity drop perceived in the middle panel of Fig. 4 (see Sec. IV.1 for more details). We have also revealed the existence of other “excited magic angles”, like $\theta\simeq 0.52,0.66^{\circ}$, where $v_{*}^{1}$ vanishes, although this degeneration phenomenon was not observed and therefore they do not manifest in the zero-mode effective Fermi velocity (12). The possible relevance and physical significance of these excited magic angles is yet to be determined.
@@
Figure 5: Energies of quasi-zero $\psi_{0}^{\pm}$ (thickest curves) and first excited $\psi_{1}^{\pm}$ Hamiltonian eigenstates [conduction (+) in red and valence (-) in blue] as a function of the twist angle $\theta$ (in degrees). Energies $E_{0}^{\pm}$ and $E_{1}^{\pm}$ degenerate at the twist angle $\theta\simeq 0.4^{\circ}$ (marked with a grey vertical gridline). This degeneracy seems to be responsible for the sudden drop of fidelity and the outlier value of the Fermi velocity (12) at $\theta\simeq 0.4^{\circ}$, between the magic angles $\theta_{2}$ and $\theta_{3}$ in Fig. 4. The inset shows that $\theta\simeq 0.4^{\circ}$ actually corresponds to an excited magic angle where the Fermi velocity expectation value $v_{*}^{1}$ in Eq. (13) vanishes for the first excited conduction state $\psi_{1}^{+}$.
@@ IV.1 Fidelity-susceptibility
In Figure 4 (middle panel) we plot the fidelity $F_{\psi_{0}^{\pm}}$ of the lowest energy (quasi-zero) $E_{0}^{\pm}$ Hamiltonian eigenstates $\psi_{0}^{\pm}$ at the Dirac point $\bm{q}=\bm{0}$ as a function of $\theta\in[0.18,1.3]$ in degrees (we take a step $\delta\theta=0.0023^{\circ}$). We identify sudden drops of fidelity exactly at the magic angles $\theta_{1}=1.05^{\circ}$ (principal), $\theta_{2}=0.495^{\circ},\theta_{3}=0.35^{\circ},\theta_{4}=0.235^{\circ}$ and $\theta_{5}=0.193^{\circ}$ where the effective Fermi velocity $v_{*}$ drops to zero.
Additionally, fidelity also detects the “excited magic angle” $\theta\simeq 0.4^{\circ}$, where the effective Fermi velocity $v_{*}^{1}$ in Eq. (13) vanishes. As already commented at the end of Sec. III, the origin of this fidelity drop is linked to the fact that the quasi-zero energy $\psi_{0}^{\pm}$ and first excited $\psi_{1}^{\pm}$ Hamiltonian eigenstates almost degenerate at $\theta\simeq 0.4^{\circ}$. Fidelity is sensitive to this degeneracy since $\psi_{0}^{\pm}$ and $\psi_{1}^{\pm}$ are orthogonal, as are $\psi_{0}^{+}$ and $\psi_{0}^{-}$.
For completeness, Figure 4 (bottom panel) shows how magic angles are also detected by sudden rises of the overlap
@@ IV.2.2 RDM and entanglement at magic angles
Figure 8: Linear entropy and its derivative for the lowest-energy valence (similar results for conduction) eigenstate $\psi_{0}^{-}$ as a function of the twist angle. Magic angles are marked with vertical gray grid lines.
@@ V Conclusions
We have contributed to unraveling the magic behind the electronic properties and exotic quantum states of TBG, which make it a promising platform for future device applications in electronics, optoelectronics, and quantum computing. Tools from the field of information theory offer a framework for examining and comprehending the complex interaction of degrees of freedom in TBG. Integrating these tools with quantum many-body physics and sophisticated computational techniques offers a promising strategy for investigating the rich landscape of quantum states in TBG and topological materials in general [27, 28].
@@
We have reviewed the flattening of low-energy Moiré Bloch bands and the formation of sharp picks in the DOS at magic angles $\theta\simeq 1.05^{\circ},0.495^{\circ},0.35^{\circ},0.235^{\circ}$ and $0.193^{\circ}$. This flattening is accompanied by a vanishing of the effective Dirac-point Fermi velocity for zero energy Hamiltonian eigenstates $\psi_{0}(\theta)$ at those magic angles. Zero modes undergo critical structural changes at magic angles, and these abrupt changes are clearly detected by fidelity measurements. Fidelity also captures an anomalous behavior at $\theta\simeq 0.4^{\circ}$, where the energies of the zero modes $\psi_{0}(\theta)$ and first excited $\psi_{1}(\theta)$ states degenerate, thus causing a spurious value of the Fermi velocity due to a numerical artifact. Sharp oscillations (like “resonances”) at magic angles are also noticeable (specially at the principal one $\theta_{1}\simeq 1.05^{\circ}$) in the entanglement entropy and the interlayer hopping probabilities (off-diagonal component of the RDM to the layer sector) as a function of the the twist angle. These resonances are better detected by the corresponding derivatives. Entanglement entropy is almost maximum in the small angle region where magic angles fall, and it decreases to zero for larger values of $\theta$. Low-energy electrons seem to prefer one layer over the other for large $\theta$, as seen in the layer probabilities $P_{\ell}$, which oscillates around $P_{\ell}=0.5$ in the small angle region where magic angles fall. The interlayer hopping probability $T$ shows high variability for small angles and goes to zero for larger values of $\theta$.
High tunability with $\theta$ and manipulation with external electric and magnetic fields in general 2D materials allows for the exploration of novel quantum phenomena termed as “twistronic” [52, 9]. We believe that quantum information viewpoints can contribute to a better understanding for the physics behind and to the design of new experiments. Sudden transitions or abrupt changes in entanglement measurements, like those evidenced in this article at magic angles, can provide a wealth of opportunities for advancing quantum information devices and opens pathways to more robust, adaptable, and efficient quantum technologies. Monitoring such variations can provide a diagnostic for fault-tolerant quantum error correction protocols. Also, sudden entanglement transitions often correspond to critical points in a quantum system (e.g., phase transitions); these critical points can be highly sensitive to external perturbations, making them ideal for quantum sensing applications and devices to achieve higher precision in measuring physical quantities like magnetic fields, gravitational waves, time intervals, etc. Also to enhance the performance and reliability of quantum communication systems and secure authentication in quantum cryptographic protocols. Bilayer graphene quantum dots have proved to reliably encode quantum information in the valley degree of freedom [53], making it a promising quantum computation platform for electrically controllable robust qubits. Tunability with $\theta$ can offer here new possibilities since we have seen that interlayer hopping probability $T$ and entanglement entropy show noticeable resonances at magic angles (especially at $\theta_{1}\simeq 1.05^{\circ}$).
@@
Novel quantum phenomena and new technological possibilities are also expected when adding a stacking degree of freedom and going from bilayer to multilayer arrangements. For example, twisted trilayer graphene seems to provide a promising platform for experimental realization of correlated topological states [54, 55]. Rhombohedral pentalayer graphene provides an ideal platform for exploring charge fractionalization and (non-Abelian) anyonic braiding at zero magnetic field [56]. Multilayer arrangements also increase the number of internal quantum degrees of freedom to the electron necessary to encode logical qubits. Therefore, investigations in this direction are worth to explore.
Note added: during the period of review of this article, we have been informed of the existence of calculations based on atomistic tight-binding models [57, 58], including atomic relaxation in the TBG model, where second (and higher) magic twist angles smaller than the principal one $\theta_{1}\simeq 1.1^{\circ}$ are not predicted. Atomic reconstruction is negligible for large twist angles but it turns out to be essential for small ones since it significantly modifies the stacking structure of TBGs, compared to the ideal isolated graphene lattices (see e.g. [59, 60] for experiments). This fact deserves a separate study. Our main goal in this article has been to provide model-independent information-theoretic measures to capture all magic angles whenever they are present in a particular TBG Hamiltonian model.